OpenStudy (yureiotaku):

The graph of a quadratic function has a domain of (-∞, ∞) and range of [4, ∞). In two or more complete sentences, explain how the given range of the function can help you to determine whether the graph opens up or down. I just really know how to do it, I have studied online and can't find anything on it.

OpenStudy (yureiotaku):

I really want to know*

OpenStudy (danjs):

The range is the set of values for Y on the X-Y plane, for y=f(x)

OpenStudy (danjs):

4 to infinity, must mean it opens "up" towards larger positive Y values

OpenStudy (yureiotaku):

@DanJS Thank you, thank you for the explanation also, I had a few questions like that but I was stuck on how to figure it out!!!

OpenStudy (danjs):

welcome

OpenStudy (danjs):

remember it is different if you have x as a function of y x = f(y)